6. The USD is strong comparing to EUR now. A EU company afraid the appreciation pressure on their USD investment. The manager comes to the bank that you work for to hedge the $3m investment due in 6 months. The following information shows up on your screen:

Spot exchange rate: $1.1230/€ - $1.1265/€ US interest rates are 2% for investing and 4.5% for borrowing.

EU interest rates are 0.1% for investing and 2.5% for borrowing.

Please remember: Interest rates are quoted as annualized rate.

The bank that you work for demands a 0.85% profit margin for all forward contract. How much is the forward quotation that you will provide to your corporate customer?
Answer #1

The Eu company wants to invest $3m so we select US interest rate for investing as rate of home country. Accordingly it will borrow in EU, so EU borrowing rates will be selected as rate of foreign company.

The quote is an indirect quote, so for an investment of $3m, the spot rate applicable for EU company will be $1.1230/€

Forward rate = Spot rate * [(1 + Rh) / (1 + Rf)]

= 1.1230 * [(1 + 0.02/2) / (1 + 0.025/2)]

= 1.1230 * (1.01 / 1.0125)

= 1.1202

Bank quote = Forward rate + Premium

= 1.1202 + 0.85% of 1.1202

= 1.202 + 0.0095

= $1.1297/€